# A Multi-Messenger Search for the Supermassive Black Hole Binary in 3C 66B with the Parkes Pulsar Timing Array

### v1

@@ Appendix B Analysis of the 3-mm data
Next, $(\beta,\iota)$ are required in the calculation of the absolute value of the rotational velocity of one of the binary components, which has a jet, around the center of mass. This velocity causes a Doppler boosting of the flux. We solve Equation 4 from Iguchi et al. (2010) for this velocity, obtaining
@@
Considering a Keplerian motion in a circular orbit, we use the following expressions of the orbital radius $R$ and the orbital separation $a$,
@@
Introducing $x\equiv R/a$, one finds that $x=m/M_{\text{tot}}$ with $m$ the mass of a lighter black hole for the case where a heavier black hole emits a jet (smaller radius). Similarly, one finds $x=M/M_{\text{tot}}$ with $M$ the mass of a heavier black hole for the case where a lighter black hole emits a jet (larger radius). Apparently, for the calculation of the chirp mass, these two cases reduce to the same answer due to a symmetry in x(1x)x(1-x) in the final expression
@@ Appendix C Pulsar Distances
Pulsar distances are required in the calculation of the expected signal when modeling the pulsar term of the CGW signal. Pulsar distances are measured using very-long-baseline-interferometry (VLBI) astrometry, optical astrometry, timing, and the Shklovskii (1970) effect. A list of the priors used for each pulsar, as well as the technique used to find this distance, is listed in Table 2. We impose a Gaussian prior on pulsar distance in accordance with the estimates found in the literature. When no such estimates are available for a given pulsar, we impose a sufficiently conservative prior $\mathcal{N}(1000,200)~\text{[pc]}$. Distance estimates based on pulsar dispersion measure are not used due to the large uncertainties in the electron density model (Yao et al., 2017).
@@
J0614-3329 $1000\pm 200$ - Ellis & Van Haasteren (2017)
J0711-6830 $1000\pm 200$ - Ellis & Van Haasteren (2017)
J0900-3144 $1000\pm 200$ - Ellis & Van Haasteren (2017)
@@
J1125-6014 $1000\pm 200$ - Ellis & Van Haasteren (2017)
J1446-4701 $1000\pm 200$ - Ellis & Van Haasteren (2017)
J1545-4550 $1000\pm 200$ - Ellis & Van Haasteren (2017)
@@
J2241-5236 $1000\pm 200$ - Ellis & Van Haasteren (2017)
Table 2: Table describing the pulsar distances, the methods used to obtain them and the source from which they were obtained. If no distance was available, the method is listed as “-”.

### v2

@@ Appendix B Analysis of the 3 mm Data
Next, $(\beta,\iota)$ are required in the calculation of the absolute value of the rotational velocity of one of the binary components, which has a jet, around the center of mass. This velocity causes Doppler boosting of the flux. We solve Equation 4 from Iguchi et al. (2010) for this velocity, obtaining
@@
Considering Keplerian motion in a circular orbit, we use the following expressions of the orbital radius $R$ and the orbital separation $a$,
@@
Introducing $x\equiv R/a$, one finds that $x=m/M_{\text{tot}}$ with $m$ the mass of the lighter black hole for the case where a heavier black hole emits a jet (smaller radius). Similarly, one finds $x=M/M_{\text{tot}}$ with $M$ the mass of the heavier black hole for the case where the lighter black hole emits a jet (larger radius). The total mass, $M_{\text{tot}}$, is obtained by using the relationship between the central stellar velocity dispersion and black hole masses (Ferrarese and Merritt, 2000; Merritt and Ferrarese, 2001). Using this relationship Noel-Storr et al. (2007) obtain a total mass for 3C 66B of $1.9^{+1.0}_{-0.8}\times 10^{9}\ M_{\odot}$. Apparently, for the calculation of the chirp mass, these two cases reduce to the same answer due to a symmetry in x(1x)x(1-x) in the final expression
@@
The value of x(x1)x(x-1) is also known as the symmetric mass ratio.
@@ Appendix C Pulsar Distances
Pulsar distances are required in the calculation of the expected signal when modeling the pulsar term of the CGW signal. Pulsar distances are measured using very long baseline interferometry (VLBI) astrometry, optical astrometry, timing, and the Shklovskii (1970) effect. A list of the priors used for each pulsar, as well as the technique used to find this distance, are listed in Table 2. We impose a Gaussian prior on pulsar distance in accordance with the estimates found in the literature. When no such estimates are available for a given pulsar, we impose a sufficiently conservative prior $\mathcal{N}(1000,200)~\text{[pc]}$. Distance estimates based on pulsar dispersion measures are not used due to the large uncertainties in the electron density model (Yao et al., 2017).
Table 2: Table Describing the Pulsar Distances, the Methods Used to Obtain Them, and the Source from Which They Were Obtained.
@@
J0614-3329 $1000\pm 200$ Ellis and Van Haasteren (2017)
J0711-6830 $1000\pm 200$ Ellis and Van Haasteren (2017)
J0900-3144 $1000\pm 200$ Ellis and Van Haasteren (2017)
@@
J1125-6014 $1000\pm 200$ Ellis and Van Haasteren (2017)
J1446-4701 $1000\pm 200$ Ellis and Van Haasteren (2017)
J1545-4550 $1000\pm 200$ Ellis and Van Haasteren (2017)
@@
J2241-5236 $1000\pm 200$ Ellis and Van Haasteren (2017)
Note. If no distance was available, the method is listed as “…”.